OpenStudy (anonymous):

An automobile has 20 inch diameter wheels. If the wheels revolved three times after the brakes were applied, the stopping distance was approximately?

OpenStudy (ash2326):

Diameter of wheels 20 inch so radius = 10 inch Wheels revolved 3 times after the breaks were applied , so the distance would be 3 times the perimeter Perimeter of circle= \(2 \pi r\) so Distance covered = \(3 \times 2 \times \pi \times r\) Can you solve it now???

OpenStudy (anonymous):

I got 3 *2* 3.14*10= 188.4. But thats not one of my choices on my paper?

OpenStudy (ash2326):

What are the options???

OpenStudy (anonymous):

A. 26.2 b. 157 c. 7.85 d. 15.7

OpenStudy (ash2326):

are the units given in options??

OpenStudy (anonymous):

yes they are all ft..

OpenStudy (ash2326):

feet:D:D \[3 *2* 3.14*10= 188.4\ inches\] 1 feet = 12 inches so \[Distance in feet= \frac{188.4}{12}\]

OpenStudy (anonymous):

okay so i got 15.45 ft.

OpenStudy (ash2326):

You should recalculate it'd be 15.7 188.4/12=(180+8.4)/12=15+0.7= 15.7 feet

OpenStudy (anonymous):

okay i got it now.